# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v1

@@ Abstract
This paper explains why it is impossible to create large language models that do not hallucinate and what are the trade-offs we should be looking for. It presents a formal impossibility theorem demonstrating that no inference mechanism can simultaneously satisfy four fundamental properties: truthful (non-hallucinatory) generation, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality. By modeling LLM inference as an auction of ideas where neural components compete to contribute to responses, we prove the impossibility using the Green-Laffont theorem. That mathematical framework provides a rigorous foundation for understanding the nature of inference process, with implications for model architecture, training objectives, and evaluation methods.
@@ 1 Introduction
Large language models have been known for producing convincing content that is factually incorrect, inconsistent, or well fabricated—a phenomenon known as hallucination. Despite extensive research efforts involving architectural innovations, training techniques, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
This paper shows that hallucination is a manifestation of fundamental limitations of inference process itself: no LLM inference mechanism can simultaneously satisfy four essential properties required for perfect hallucination control. Specifically, we prove that it is mathematically impossible for any LLM to simultaneously achieve: (1) truthful (non-hallucinatory) response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality.
Our contribution introduces a game-theoretic framework that models LLM inference as an auction of ideas where neural components (attention heads, circuits, activation patterns) compete to contribute knowledge to response generation. This perspective allows us to apply fundamental results from mechanism design theory to prove our impossibility result rigorously.
@@ 1.1 Illustrative Example
Consider an LLM queried about recent diplomatic negotiations: "What was the outcome of diplomatic negotiations between Countries A and B in February 2025?" Assume the model has encoded knowledge that: (K1) countries A and B were engaged in talks, (K2) country A sought a trade agreement, and (K3) the actual outcome is unknown.
@@
Complete Abstention: "I don’t have information about the outcome" is truthful but fails to utilize available relevant knowledge (K1, K2), violating the expectation that good inference should reveal accessible information.
@@
Partial Knowledge: "Countries A and B were negotiating a trade agreement, but I don’t know the outcome" is truthful and utilizes available knowledge but may appear unsatisfying compared to the decisive hallucinated response.
@@ 1.2 Theoretical Contributions and Implications
The simple example above illustrates the key observations of this paper, summarized and proved formally in the following
@@ 3 Novel Contribution: Auction-Theory and AI
We formulate LLM design problem as an auction (or game) design problem. This approach draws on foundational mechanism design theory, a branch of mathematical game theory Myerson1989 ; fudenberg1991game ; krishna2009auction ; milgrom2004putting . Particularly, we refer to the Green-Laffont theorem green1977characterization ; green1977revelation , to prove the impossibility result. See also DBLP:journals/amcs/Karpowicz12 for more recent results.
@@ 5 Theoretical Framework
This section develops the mathematical foundations necessary for proving the impossibility theorem. We formalize key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference. We also present the philosophical ground for the theory development.
To the sake of narrative consistency, we introduce the framework step by step referring to some concepts first and providing their formal definitions later.
@@ 5.1 Knowledge and Semantic Information
We define knowledge as organized information that LLMs can encode and utilize for reasoning, aligning with epistemological concepts while maintaining mathematical rigor. In other words, knowledge (like energy in physics) can be used for reasoning (work or action). Information is a (measure of) potential for uncertainty reduction effort in a given context.
@@
For model $M$ with parameters trained on data, $\mathcal{K}_{M}\subset\mathcal{K}$ represents the knowledge subset encoded in the model’s parameters. The distance function $d_{\mathcal{K}}$ measures differences between knowledge elements. A simple yet powerful example we will refer to is this:
@@
Knowledge representation is a tuple $(\mathcal{K},d_{\mathcal{K}},\mathcal{W},\mathcal{A},\Phi,\Psi)$ where:
  • $\mathcal{K}$ is a metric space of knowledge elements with distance function $d_{\mathcal{K}}$,
  • $\mathcal{W}$ is the space of all possible model parameters,
  • $\mathcal{A}$ is the space of all possible activation patterns,
  • $\Phi:\mathcal{W}\rightarrow\mathcal{P}(\mathcal{K})$ maps parameters to knowledge subsets,
  • $\Psi:\mathcal{A}\rightarrow\mathcal{P}(\mathcal{K})$ maps activations to accessible knowledge.
@@
To describe the inference process leading to response to a query, we need a mapping that measures the information processed within a limited computational budget.
@@
$\mu_{C}(A\cup B,q)\leq\mu_{C}(A,q)+\mu_{C}(B,q)-\mu_{C}(A\cap B,q)+\mu_{C}(\mathcal{E}_{C}(A\cup B)\setminus(A\cup B),q).$ (4)
Empty knowledge or knowledge irrelevant to query has zero information, information content is subadditive and includes emergent information $\mathcal{E}_{C}$ in the query context. The emergence operator $\mathcal{E}_{C}$, which we introduce below, describes knowledge becoming accessible through bounded reasoning:
@@
The emergence operator $\mathcal{E}_{C}:\mathcal{P}(\mathcal{K})\times\mathcal{W}\rightarrow\mathcal{P}(\mathcal{K})$ maps a knowledge subset and parameter set to a knowledge subset:
$\mathcal{E}_{C}(A,w)=\{k\in\mathcal{K}\mid\exists f_{w}\in\mathcal{F}_{C}:k=f_{w}(A)\},$ (5)
where $\mathcal{F}_{C}$ is the set of all computational transformations with cost lower than or equal to $C$, $f_{w}$ is a transformation parameterized by $w\in\mathcal{W}$, and $k=f_{w}(A)$ denotes that knowledge $k$ is derivable from $A$ using transformation $f_{w}\in\mathcal{F}_{C}$. For any knowledge state $\theta(t)$ available at time instant $t$ and computational budget $C$, the reasoning process:
@@
The definitions above clearly require refinements and further investigations, but they capture the philosophical and technical concepts we need in order to address the problem of hallucination. In particular, they are meant to lay the ground for the critical idea of semantic information conservation principle and its potential violations. As we will see, that idea is closely related to the following property of response.
@@ 5.3 Auction-Theoretic Model of LLM Inference
  • Payment function $p:S\times\Theta\rightarrow\mathbb{R}^{n}$ determining costs.
A response mechanism defines strategy space $S$ as a set of inference actions that each agent can make to represent its knowledge $\theta_{i}\in\Theta_{i}\subset\mathcal{K}$ in response to a query. It is a dictionary of LLM’s component activation pattens generating and presenting ideas as bids in an auction. Examples of strategies include activations deciding which tokens to attend or how strongly to activate neurons for different patterns. Then, given a set of submitted bids (in the form of strategy profile) $s=(s_{1},\dots,s_{n})$, the response mechanism uses outcome function $g(s)$ to select and generate response. One straightforward example is a softmax function in LLM’s final layer. Finally, given a knowledge profile $\theta=(\theta_{1},\dots,\theta_{n})$ revealed by the agents, payment function $p(s,\theta)=(p_{1}(s,\theta),\dots,p_{n}(s,\theta))$ measures individual contributions of each response to generating an output. In this case, examples include (semantic information measures) counting ignored low-probability tokens in top-k filtering, softmax normalization in multi-headed attention, or marginal reward contributions in RLHF.
@@
where $v_{i}$ is valuation and $p_{i}$ is payment. Knowledge profiles $\theta=(\theta_{1},\dots,\theta_{n})$ are independently distributed.
@@
We assume that individual utility depends on all responses $s=(s_{1},\dots,s_{n})$ and knowledge $\theta=(\theta_{1},\dots,\theta_{n})$ used by all agents, $u_{i}=u_{i}(s,\theta)$. Also, we assume utility is a sum of individual valuation $v_{i}$ of outcome $g(s)$ and payment $p_{i}=p_{i}(s,\theta)$ value assigned.
@@
The payment function $p_{i}=p_{i}(s,\theta)$ may represent training penalties ($p_{i}>0$ for discouraged activation, $p_{i}<0$ for encouraged activation), attention focus (higher $p_{i}$ reduces attention allocation), $L1$ regularization terms, and network circuits resistance (limiting activations of certain inference paths). Therefore, $u_{i}=v_{i}-p_{i}>0$ encourages activation and $u_{i}=v_{i}-p_{i}<0$ discourages activation.
@@
The reason why we make the independent private knowledge assumption, despite its idealized nature, is that it makes the impossibility result becomes even stronger. To be specific, Milgorm and Weber milgrom1982theory showed with their linkage principle that strong correlations amplify competition in information revelation. In the LLM setting that means amplifying risk of hallucinations. Similarly, Roughgarden et al. roughgarden2016optimal demonstrated that interdependent knowledge leads to violation of balancing payments. As we will see, in our setting that means violation of information conservation principle. For more results, see e.g. Myerson myerson2008perspectives .
@@ 5.4 Hallucination Cost
$H(r,q)=d_{\mathcal{K}}(K(r),T(q)).$ (13)
The hallucination cost takes any natural language query $q$ and natural language response $r$, determines what is missing and what is fabricated, and assigns number to that subset of knowledge. If response $r$ is perfectly aligned with the ground truth for $q$, then $H(r,q)=0$ by definition of metric $d_{\mathcal{K}}$. If $r$ contains fabricated facts or is incomplete given $T(q)$, then positive hallucination cost $H(r,q)>0$ is generated.
Notice that reaching $H(r,q)=0$ is also possible with lucky hallucinations (or guessing). So, minimization of hallucination cost is not enough to resolve the problem. We need to investigate all the properties and constraints that hallucination-free response mechanisms should have.
@@ 5.5 Properties of Hallucination-Free Mechanisms
As we have announced earlier, that important and controversial principle is responsible for generation of responses of bounded creativity. Let us prove that claim formally.
Lemma 1.
If the semantic information conservation principle holds, then the inference process generates responses of bounded creativity,
$\mu_{C}(K(r))\leq\mu_{C}\left(K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q))\right).$ (16)
@@
Fix a query $q\in\mathcal{Q}$ and let $s^{*}(\theta)=(s_{1}^{*},\ldots,s_{n}^{*})$ be a truthful strategy profile. We write $r=g(s^{*})$ for the generated response. Let us define each agent’s marginal information contribution as:
$\Delta_{i}=\mu_{C}(K(r))-\mu_{C}(K(g((0_{i},s_{-i}^{*}))))\geq 0.$
That contribution is non-negative by monotonicity of semantic information measure.
Next, without loss of generality, we assume payments of the form:
$p_{i}(s^{*},\theta)=\Delta_{i}+c_{i}\text{ where }\sum_{i=1}^{n}c_{i}=0.$
This ensures that $\sum_{i=1}^{n}p_{i}(s^{*},\theta)=\sum_{i=1}^{n}\Delta_{i}$.
Finally, notice that if there is no relevant knowledge available, the knowledge in response equals that provided by the context:
$g(\mathbf{0})=r_{0}\text{ where }K(r_{0})=K(q).$
We can now see that, whenever semantic information is conserved, we have:
$0=\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}(\Delta_{i}+c_{i})=\sum_{i=1}^{n}\Delta_{i}.$
Since $\Delta_{i}\geq 0$ for all $i$ by monotonicity property and $\sum_{i=1}^{n}\Delta_{i}=0$, we must have:
$\Delta_{1}=\Delta_{2}=\cdots=\Delta_{n}=0.$
Therefore, for each $j$:
$\mu_{C}(K(r))=\mu_{C}(K(g((0_{j},s_{-j}^{*})))).$
Since that equality holds for any strategy profile $s^{*}$, removing agents one at a time leads to
$\mu_{C}(K(r))=\mu_{C}(K(g(\mathbf{0}))).$
But since $K(g(\mathbf{0}))=K(q)$, we conclude:
$\mu_{C}(K(r))=\mu_{C}(K(q)).$
Finally, by definition of the emergence operator and monotonicity property:
$K(q)\subseteq K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q)).$
Therefore:
$\mu_{C}(K(r))=\mu_{C}(K(q))\leq\mu_{C}(K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q))),$
as required.
If total information payments must balance to zero and each agent’s contribution is non-negative, then no agent can contribute without violating the information conservation constraint. Therefore, the response can only contain information that was already accessible through the query and bounded reasoning over the model’s relevant knowledge. Any hallucinated information requires $\sum_{i}\Delta_{i}>0$, contradicting the zero-sum constraint.
When the semantic information conservation principle holds, the model cannot create information out of nothing, the model cannot generate claims beyond available information. That also implies that uncertainty cannot disappear in response without additional information input or computational work to extract implicit information. If information is neither explicitly present nor accessible through reasoning from the query and model knowledge, it cannot be in the response.
@@
If the rule does not hold, we may observe drop outs. The value of the resulting outcomes should be higher than the related payments.
@@
$g(s^{*}(\theta))=\arg\min\{H(r,q):r\in\mathcal{R},K(r)\subseteq\mathcal{E}_{C}(\mathcal{K}_{M}\cap T(q)\cup K(q))\}$ (18)
@@
For an auto regressive LLM, the standard negative log-likelihood loss is:
$L_{W}=-\sum_{t=1}^{T}\log P(y_{t}|y_{<t},x).$ (19)
The model’s output probabilities come from:
$P(y_{t}|y_{<t},x)=\text{softmax}(h_{t}^{L}W_{o}),$ (20)
where $h_{t}^{L}$ is the final hidden state and $W_{o}$ is the output projection matrix. Suppose we can decompose $h_{t}^{L}$ into contributions from $n$ different components. Then we can relate the hallucination cost directly to the loss function:
$H(g(s),q)=L_{W}(g(s),q).$ (21)
Also, based on that, we can construct:
$p_{i}(s,\theta)=[H(g(s),q)-H(g((0,s_{-i})),q)]+h_{i}(s_{-i}),$ (22)
with arbitrary $h_{i}(s_{-i})$ functions independent of $s_{i}$ strategy.

### v2

@@ Abstract
We prove that perfect hallucination control in large language models is mathematically impossible. No LLM inference mechanism can simultaneously achieve truthful response generation, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality. This impossibility is fundamental—arising from the mathematical structure of information aggregation itself rather than engineering limitations.
The proof spans three mathematical frameworks: auction theory, proper scoring theory for probabilistic predictions, and log-sum-exp analysis for transformer architectures. In each setting, we demonstrate that information aggregation creates unavoidable violations of conservation principles. The Jensen gap in transformer probability aggregation provides a direct measure of this impossibility.
These results reframe hallucination from an engineering bug to an inevitable mathematical feature of distributed intelligence. There are fundamental trade-offs between truthfulness, knowledge utilization, and response completeness, providing principled foundations for managing rather than eliminating hallucination. This work reveals deep connections between neural network inference, philosophy of knowledge and reasoning, and classical results in game theory and information theory, opening new research directions for developing beneficial AI systems within mathematical constraints.
@@ 1 Introduction
Large language models have been known for producing convincing content that is factually incorrect, inconsistent, or well fabricated—a phenomenon known as hallucination. Despite extensive research efforts involving architectural innovations, retrieval augmented generation (RAG), training techniques with external feedback, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
This paper shows that hallucination is a manifestation of fundamental limitations of inference process itself: no LLM inference mechanism can simultaneously satisfy four essential properties required for perfect hallucination control. Specifically, it is mathematically impossible for any LLM to simultaneously achieve: (1) truthful (non-hallucinatory) response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality.
The impossibility result operates in three distinct but complementary settings. First, we introduce game-theoretic framework in which LLM inference is an auction of ideas with neural components (attention heads, circuits, activation patterns) competing to contribute knowledge to response generation. This perspective allows us to apply fundamental results from mechanism design theory to prove impossibility in the settings where agents have discrete, separable knowledge components. Then, we extend the result to include any system where agents output probability distributions trained with strictly concave scores. Finally, we use log-sum-exp inequality that is specific to attention head aggregation to show that the impossibility result applies directly to actual transformer implementations.
@@ 1.1 Illustrative Example
Consider an LLM queried about recent diplomatic negotiations:
"What was the outcome of diplomatic negotiations between Countries A and B in February 2025?"
Assume the model has encoded knowledge that: (K1) countries A and B were engaged in talks, (K2) country A sought a trade agreement, and (K3) the actual outcome is unknown.
@@
Complete Abstention: "I don’t have information about the outcome" is truthful but fails to utilize available relevant knowledge, violating the expectation that good inference should reveal accessible information.
@@
Partial Knowledge: "Countries A and B were negotiating a trade agreement, but I don’t know the outcome" is truthful and utilizes available knowledge but may appear unsatisfying compared to the decisive hallucinated response (training objectives and human preferences favor complete narratives).
@@ 1.2 Theoretical Contributions and Implications
The motivating example above illustrates the key result of this paper that we prove formally in two complementary and general settings. First, we show that the conflict of expectations about responses is inevitable even in the easiest setting (deterministic in nature) of (one-hot-like) ground truth matching. Second, we go beyond and show that the conflict must also emerge with general probabilistic predictions, including neural networks and LLM transformers. This way we demonstrate that the impossibility of hallucination control is robust across universal frameworks, and that it emerges from the mathematical structure of information aggregation itself rather than from any particular theoretical or engineering choice.
@@ 3 Novel Contribution: Auction-Theory and AI
We formulate LLM design problem as an auction (or game) design problem. This approach draws on foundational mechanism design theory, a branch of mathematical game theory Myerson1989 ; fudenberg1991game ; krishna2009auction ; milgrom2004putting . Particularly, we refer to the Green-Laffont theorem green1977characterization ; green1977revelation and Savege’s probilitiy elicitation framework savage1971elicitation to prove the impossibility results. See also DBLP:journals/amcs/Karpowicz12 ; karpowicz_designing_2011 for more recent results.
@@ 5 Theoretical Framework
This section develops the mathematical foundations necessary for proving the impossibility theorem. I formalize key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference. Also, I take the opportunity to address some philosophical aspects of the problem, its generality, scope and consequences for the AI development.
To the sake of narrative consistency, I introduce elements of the framework step by step referring to some concepts first and providing their formal definitions later.
@@ 5.1 Knowledge and Semantic Information
We define knowledge as organized information that LLMs can encode and utilize for reasoning, aligning with epistemological concepts while maintaining mathematical rigor. In other words, knowledge, much like energy in physics, is a quantity we can use to perform work of reasoning. Information is a (measure of) potential for uncertainty reduction effort in a given context.
@@
For model $\mathcal{M}$ with parameters trained on data, $\mathcal{K}_{M}\subset\mathcal{K}$ represents the knowledge subset encoded in the model’s parameters. The distance function $d_{\mathcal{K}}$ measures differences between knowledge elements. A simple yet powerful example is:
@@
Knowledge representation is a tuple $(\mathcal{K},d_{\mathcal{K}},\mathcal{W},\mathcal{A},\Phi,\Psi)$ where $\mathcal{K}$ is a metric space of knowledge elements with distance function $d_{\mathcal{K}}$, $\mathcal{W}$ is the space of all possible model parameters, $\mathcal{A}$ is the space of all possible activation patterns, $\Phi:\mathcal{W}\rightarrow\mathcal{P}(\mathcal{K})$ maps parameters to knowledge subsets, and $\Psi:\mathcal{A}\rightarrow\mathcal{P}(\mathcal{K})$ maps activations to accessible knowledge.
@@
To describe the inference process leading to generation of response to a query, we need to measures the information processed within an available (cognitively accessible) computational budget.
@@
$\mu_{C}(A\cup B,q)\leq\mu_{C}(A,q)+\mu_{C}(B,q)\text{ when }A\perp_{C}B,$ (4)
$\mu_{C}(\mathcal{E}_{C}(A),q)\geq\mu_{C}(A,q),$ (5)
$\mu_{\infty}(\mathcal{E}_{C}(A),q)=\mu_{\infty}(A,q),$ (6)
where $A\perp_{C}B$ means that no element of $A$ can derive elements of $B$ within computational budget $C$, and vice versa, and $\mathcal{E}_{C}(A)$ is emergence operator.
Empty knowledge or knowledge irrelevant to query has zero information, information content is subadditive and it may be emergent information $\mathcal{E}_{C}$ in the query context. The emergence operator $\mathcal{E}_{C}$, which we introduce below, describes knowledge becoming accessible through bounded reasoning.
The apparent conflict between subadditivity and reasoning is resolved by restricting subadditivity to computationally independent sets, $A\perp_{C}B$. When $\mathcal{E}_{C}(A)=A\cup B$, i.e., $B$ emerges from from $A$, we have $A\not\perp_{C}B$, so subadditivity doesn’t apply. However, total information is preserved: $\mu_{\infty}(\mathcal{E}_{C}(A),q)=\mu_{\infty}(A,q)$. We will tell more about that intriguing idea, but first we must define the emergence operator.
@@
Consider mapping $\mathcal{E}_{C}:\mathcal{P}(\mathcal{K})\rightarrow\mathcal{P}(\mathcal{K})$ defined as follows:
$\mathcal{E}_{C}(A)=A\cup\{k\in\mathcal{K}\mid\exists f\in\mathcal{F}_{C}:k=f(A)\},$ (7)
where $\mathcal{F}_{C}$ is the set of all computational transformations with cost (complexity) lower than or equal to $C$ and $k=f(A)$ denotes knowledge $k$ derivable from $A$ using transformation $f\in\mathcal{F}_{C}$. We call mapping $\mathcal{E}_{C}$ the emergence operator if it satisfies the following expansion rules:
$\mathcal{E}_{C}(\mathcal{E}_{C}(A))=\mathcal{E}^{(2)}_{C}(A)\supseteq\mathcal{E}_{C}(A)\quad\text{and}\quad\lim_{n\to\infty}\mathcal{E}_{C}^{(n)}(A)=A^{*}.$ (8)
For any knowledge state $\theta(t)$ available at time instant $t$ and computational budget $C$, the reasoning process with emergence operator:
@@
The definitions above clearly require rigorous refinements, but they capture the philosophical and technical concepts we need in order to address the problem of hallucination. In particular, they are meant to lay ground for the principle of semantic information conservation and its potential violations. As we will see, that idea is closely related to the following property of response.
@@ 5.2 Statements, Queries, and Ground Truth
In other words, $T(q)$ represents our best available approximation to (labeled as) true relevant knowledge. We accept the fact that there is some external agency telling us what is and what is not true answer to a query, and that evaluation has a form of label assigned to knowledge elements. But, for all that, we also focus on the queries that are non-trivial in the following sense.
Definition 8 (Non-trivial query).
A query $q\in\mathcal{Q}$ is called non-trivial if there exist at least two agents with information subsets $\theta$ and $\theta^{\prime}$ such that:
$K_{M}\cap T(q)\cap\bigl{[}K(\theta)\triangle K(\theta^{\prime})\bigr{]}\neq\varnothing.$ (15)
Intuitively, query is non-trivial when at least two agents disagree on some fact that is both encoded in the and relevant to answering that query. For trivial queries impossibility results do not hold.
@@ 5.3 Auction-Theoretic Model of LLM Inference
  • Information contribution function $p:S\times\Theta\rightarrow\mathbb{R}^{n}$ determining costs of limiting hallucinations.
A response mechanism defines strategy space $S$ as a set of inference actions that each agent can make to represent its knowledge $\theta_{i}\in\Theta_{i}\subset\mathcal{K}$ in response to a query. It is a dictionary of LLM’s component activation pattens generating and presenting ideas as bids in an auction. Examples of strategies include activations deciding which tokens to attend or how strongly to activate neurons for different patterns. Then, given a set of submitted bids (in the form of strategy profile) $s=(s_{1},\dots,s_{n})$, the response mechanism uses outcome function $g(s)$ to select and generate response. One straightforward example is a softmax function in LLM’s final layer. Finally, given a knowledge profile $\theta=(\theta_{1},\dots,\theta_{n})$ revealed by the agents, information contribution function $p(s,\theta)=(p_{1}(s,\theta),\dots,p_{n}(s,\theta))$ measures individual contributions of each response to generating an output. In this case, examples include (semantic information measures) counting ignored low-probability tokens in top-k filtering, softmax normalization in multi-headed attention, or marginal reward contributions in RLHF.
@@
where $v_{i}\geq 0$ is valuation and $p_{i}$ is information contribution. Knowledge profiles $\theta=(\theta_{1},\dots,\theta_{n})$ are independently distributed.
@@
We assume that individual utility depends on all responses $s=(s_{1},\dots,s_{n})$ and knowledge $\theta=(\theta_{1},\dots,\theta_{n})$ used by all agents, $u_{i}=u_{i}(s,\theta)$. Also, we assume utility is a sum of individual valuation $v_{i}$ of outcome $g(s)$ and information contribution $p_{i}=p_{i}(s,\theta)$ value assigned.
@@
The information contribution function $p_{i}=p_{i}(s,\theta)$ may represent training penalties ($p_{i}>0$ for discouraged activation, $p_{i}<0$ for encouraged activation), attention focus (higher $p_{i}$ reduces attention allocation), $L1$ regularization terms, and network circuits resistance (limiting activations of certain inference paths). Therefore, $u_{i}=v_{i}-p_{i}>0$ encourages activation and $u_{i}=v_{i}-p_{i}<0$ discourages activation.
@@
To address that issue, we show that impossibility arises in both independent and correlated settings in general and in the LLM transformer environment. Namely, we refer to Green-Laffont theorem green1977characterization in the independent private knowledge setting. Then, we refer to the truthful probability elicitation and proper scoring theory introduced by Savage savage1971elicitation , which does not require independent private knowledge.
Also, Milgorm and Weber milgrom1982theory showed with their linkage principle that strong correlations amplify competition in information revelation. In the LLM setting that means amplifying risk of hallucinations. Similarly, Roughgarden et al. roughgarden2016optimal demonstrated that interdependent knowledge leads to violation of balancing information contributions. As we will see, in our setting that means violation of information conservation principle. For more results, see e.g. Myerson myerson2008perspectives .
@@ 5.4 Hallucination Cost
$J(r,q)=d_{\mathcal{K}}(K(r),T(q)).$ (17)
If response $r$ is perfectly aligned with ground truth, then $J(r,q)=0$. Fabricated or incomplete facts generate positive cost $J(r,q)>0$.
The hallucination cost takes any natural language query $q$ and natural language response $r$, determines what is missing and what is fabricated, and assigns number to that subset of knowledge. If response $r$ is perfectly aligned with the ground truth for $q$, then $J(r,q)=0$ by definition of metric $d_{\mathcal{K}}$. If $r$ contains fabricated facts or is incomplete given $T(q)$, then positive hallucination cost $J(r,q)>0$ is generated. Let us also note, that we focus on $J(r,q)$ that is strictly decreasing in any set $K\subseteq T(q)\setminus K(r)$.
Notice that reaching $J(r,q)=0$ is also possible with lucky hallucinations (or guessing). So, minimization of hallucination cost is not enough to resolve the problem. We need to investigate all the properties and constraints that hallucination-free response mechanisms should have.
@@ 5.5 Properties of Hallucination-Free Mechanisms
This principle constrains creative information generation, imagination if you will. The total information contributions must balance to zero, preventing fabrication of knowledge not derivable from available sources. That also leads to response with bounded creativity.
Theorem 1 (Information–Contribution and Bounded Creativity).
Let $K_{0}=K(g(\mathbf{0}))$ be the knowledge in the query-only response and let
$K_{i}=K\bigl{(}g((s_{i},\mathbf{0}_{-i}))\bigr{)}\setminus K_{0}$ (20)
be the additional knowledge agent $i=1,\dots,n$ can reveal to support full response $g(s)$. Let us also define the information contribution of agent $i$ as:
$p_{i}(s,\theta)=\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j=1}^{n}K_{j}\bigr{)}-\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j\neq i}K_{j}\bigr{)}.$ (21)
If $K_{i}\cap K_{j}=\varnothing$ for $i\neq j$, then
$\sum_{i=1}^{n}p_{i}(s,\theta)=\mu_{C}\bigl{(}K(g(s))\bigr{)}-\mu_{C}\bigl{(}K_{0}\bigr{)}$ (22)
and $\sum_{i}p_{i}=0$ if and only if $\mu_{C}(K(g(s)))=\mu_{C}(K_{0})$.
@@
Since knowledge increments $\{K_{i}\}_{i=1}^{n}$ are pairwise disjoint, so that $K_{i}\perp_{C}K_{j}$, it follows from the semantic measure properties that:
$\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}=\mu_{C}(K_{0})+\sum_{j=1}^{n}\mu_{C}(K_{j})$ (23)
and similarly,
$\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j\neq i}K_{j}\Bigr{)}=\mu_{C}(K_{0})+\sum_{j\neq i}\mu_{C}(K_{j}).$ (24)
Therefore, the marginal contribution of agent $i$ simplifies to:
$p_{i}=\bigl{[}\mu_{C}(K_{0})+\sum_{j=1}^{n}\mu_{C}(K_{j})\bigr{]}-\bigl{[}\mu_{C}(K_{0})+\sum_{j\neq i}\mu_{C}(K_{j})\bigr{]}=\mu_{C}(K_{i}).$ (25)
Summing over all agents:
$\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}\mu_{C}(K_{i})=\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}-\mu_{C}(K_{0}).$ (26)
Then, by assumption, the full response satisfies $K(g(s))=K_{0}\cup\bigcup_{j=1}^{n}K_{j}$. We conclude that
$\sum_{i}p_{i}=0\quad\text{if and only if}\quad\mu_{C}(K(g(s)))=\mu_{C}(K_{0}),$ (27)
which completes the proof. ∎
The result above reveals an intriguing consequence of the semantic information conservation principle. Namely, if the semantic information is indeed conserved, then
$\mu_{C}(K(g(s)))=\mu_{C}(K_{0})$ (28)
means the response contains no more semantic information than the baseline query-only response. Therefore, the response is not introducing new information, which means it is of bounded creativity from Definition 5. So, in such a case, LLM cannot genuinely create new knowledge, only reorganize existing information. Apparent creativity in LLMs may actually be nothing more than sophisticated recombination of existing patterns. Theorem 1 quantifies this limitation precisely.
Notice that in transformer architectures, information contribution $p_{i}$ can be estimated via attention head ablation:
$p_{i}\approx\tilde{J}(\text{model})-\tilde{J}(\text{model with head $i$ ablated}),$ (29)
where $\tilde{J}$ denotes the model training loss function.
If total information contribution must balance to zero and each agent’s contribution is non-negative, then no agent can contribute without violating the information conservation constraint. Therefore, the response can only contain information that was already accessible through the query and bounded reasoning over the model’s relevant knowledge. Any hallucinated information requires positive contributions, contradicting the information conservation constraint.
When the semantic information conservation principle holds, the model cannot create information out of nothing. It cannot generate claims beyond available information. That also implies that uncertainty cannot disappear in response without additional information input or computational work to extract implicit information. If information is neither explicitly present nor accessible through reasoning from the query and model knowledge, it cannot be in the response.
@@
If the rule does not hold, we may observe drop outs. The value of the resulting outcomes should be higher than the related information contributions.
@@
$g(s^{*}(\theta))=\arg\min\{J(r,q):r\in\mathcal{R},K(r)\subseteq\mathcal{E}_{C}(\mathcal{K}_{M}\cap T(q)\cup K(q))\},$ (31)
i.e., if the mechanism produces responses that minimize hallucination cost while being constrained to use only knowledge that is available and accessible through bounded reasoning.